[ I have an 85 for my semester grade, what grade do I need to make on my final to make an A (90%) in the class? My final is 10% of the grade. ]
Answer by TimothyLamb(4379)

---
90*85 + 10f = 100*90
10f = 100*90 - 90*85
f = (100*90 - 90*85)/10
f = 135
---
you need 135% on the final, which isn't possible, so you cannot get 90% in the class.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php